# Transcendence: Generative Models Can Outperform The Experts That Train Them

### v1

@@ Abstract
Generative models are trained with the simple objective of imitating the conditional probability distribution induced by the data they are trained on. Therefore, when trained on data generated by humans, we may not expect the artificial model to outperform the humans on their original objectives. In this work, we study the phenomenon of transcendence: when a generative model achieves capabilities that surpass the abilities of the experts generating its data. We demonstrate transcendence by training an autoregressive transformer to play chess from game transcripts, and show that the trained model can sometimes achieve better performance than all players in the dataset. [^1] To play with our models, code, and data, please see our website at https://transcendence.eddie.win.  We theoretically prove that transcendence is enabled by low-temperature sampling, and rigorously assess this experimentally. Finally, we discuss other sources of transcendence, laying the groundwork for future investigation of this phenomenon in a broader setting.
@@ 1 Introduction
Our objective is to formalize the notion of transcendence and focus narrowly on this source of improvement over the experts: the removal of diverse human biases and errors. We prove that this form of denoising is enabled by low-temperature sampling, which implicitly induces a majority vote. This draws a subtle but deep connection from our new setting to a rich prior literature on model ensembling [1, 6, 19], enabling several key results. We precisely characterize the conditions under which transcendence is possible, and give a rigorous theoretical framework for enabling future study into the phenomenon. To test the predictive power of our theory, we then empirically demonstrate these effects. Digging deeper into the effects of majority voting, we show that its advantage is primarily due to performing much better on a small subset of states—that is, under conditions that are likely key to determining the outcome of the game. We also find that diversity in the data is a necessary condition for practically effective majority voting, confirming our theoretical findings. In short:
@@ 4.1 Experimental Setup
We trained several $50$M parameter autoregressive transformer decoders following best practices from modern large model training, including a cosine learning rate schedule and similar batch size-learning rate ratios as prescribed by the OPT-175B team [37]. Our dataset consists of human chess games from the lichess.org open source database from January 2023 to October 2023. In total, this dataset contains approximately one billion games. In this setting, an expert is a specific individual player. To test for transcendence, we truncate this dataset by a maximum rating, so that during training a model only sees data up to a given rating. We train our model on the next-token prediction objective, and represent our chess games as Portable Game Notation (PGN) strings, such as 1.e4 e5 2.Nf3 Nc6 3.Bb5... 1/2-1/2. Note that we do not give any rating or reward information during training - the only input the model sees are the moves and the outcome of the game. We tokenize our dataset at the $32$-symbol character level. (For further details, see Appendix E.) Our model plays chess “blind”—without direct access to the board state—and, furthermore, is never explicitly given the rules of the game: at no point is play constrained to valid outputs for a given piece or board state. Nontrivial chess skill is therefore not straightforward to acquire, and if not for the surprising capabilities of modern large transformers, one might imagine such a model would fail to learn even the basic rules of playing chess. This blindfolded setting has also been studied by prior work [23, 30], as discussed further in section 5.
One gap between our theory and practice is that in our theory, we assume that each expert is defined over the entire input space $\mathcal{X}$. However, in the chess setting this is extremely unlikely to be the case after around move $15$, as there are more unique chess games than atoms in the universe due to the high branching factor of the game tree. To address this gap, we visualize the latent representation of our model in Figure 3, where we find the model is able to capture meaningful semantics regarding both the relative advantage of a state, as well as the identity of the black and white player. This points to the ability of our model to generalize by compressing games into some shared latent representation, enabling experts to generalize to unseen states, bridging this gap between theory and practice.
@@ 4.2 Experimental Results
$d^{f}$ refers to the state visitation distribution [31] when following $f$ in a sequential setting - informally, this can be thought of the distribution of states seen when sampling from $f$ with a fixed transition function that takes in an input $x$, a output $y$, and outputs a next input $x$. Here, that transition function is given by the rules of chess and the opponent player. Given this favor function, we can now quantitatively explore the effects that lead to transcendence by setting the baseline $f$ to be the original imitation-learned probability distribution (temperature $\tau=1$), and $f^{\prime}$ as a low-temperature intervention on $f$ (e.g. temperature $\tau=0$). We can empirically calculate the reward by using the evaluation function [22] of Stockfish, an expert neural reward function that Stockfish uses to calculate its next move. This reward function is a neural network trained to predict the probability of winning through a sigmoid on a linear combination of handcrafted expert heuristics, such as amount of material versus opponent material, and number of moves to a potential checkmate.
@@
In Figure 4, we find that lowering the temperature has the effect of skewing the expected reward distribution to the left, especially for the green $\tau=0.001$ distribution. This implies that the model does not improve the expected reward by a small amount for many game states, but rather improves the expected reward by a relatively large amount for a few game states. Thus, $\tau=0.001$ improves the expected reward (probability of winning) by an average of $\mathbf{2.15\pm 0.17\%}$, but for some states, this expected improvement is over 5%. Note that the original temperature expected reward can be thought of as a Dirac distribution centered at 00.  This answers our research question in this section: Low-temperature sampling is able improves the expected reward by relatively large amounts for some specific game states, which is likely why the ChessFormer $1000$ and $1300$ model was able to achieve transcendence.
@@
In Table 1, we present the statistics of the favor function for different temperature values. From this table, we observe that as the temperature decreases, the top-$k$ accuracies monotonically increase, suggesting that the model becomes more consistent in selecting good moves. We also observe that although the model improves as temperature decreases, the probability of winning is still below $50\%$, meaning our model should tend to lose more games than it wins against Stockfish $1$. This is matches with our results in Figure 1, as the rating of Stockfish $1$ is also higher than the reported rating for $\tau=0.001$ ($1550$ for Stockfish 1 vs $\sim 1450$ for Chessformer $1000$). Overall, the analysis of the advantage statistics provides further evidence for the effectiveness of low-temperature sampling in inducing transcendence in chess models.
@@
As we note in subsection 3.4, our theory requires dataset diversity as a necessary condition for enabling transcendence. Importantly, we find in Figure 1 that not all models are able to transcend. Unlike ChessFormer 1000 or 1300, the Chessformer 1500 fails to transcend. We hypothesize that this is due to the fact that in the band of ratings from $1000$ to $1500$, diversity does not significantly increase. If this is true, a $1000$ rated player can be thought of as a noisy $1500$ rated player, but a $1500$ rated player cannot be thought of as a noisy $2000$ rated player. In this section we ask the following research question: Is diversity in data required for enabling transcendence?
@@
Importantly, we cannot calculate this normalized entropy for every state, as most states after move $16$ in the midgame and before the engame are unique within the dataset and we therefore observe just a single action for thus states. Therefore our metric is limited in that it only considers opening moves, the beginning of the midgame, and the endgame. We consider only common states with greater than $100$ actions by sampling $1,000,000$ games from each dataset. The average entropy confirm our hypothesis: The $<1500$ cut off dataset has on average less diversity than the $<1300$ dataset, which has is again less than the $<1000$ dataset. This points towards answering our research question in the affirmative; Chessformer $1500$ likely is not transcendent due to a lack of diversity in its dataset. If the entropy stayed constant for each dataset, this would imply a similar level of diversity for each. In such a case, we would expect that ChessFormer $1500$ likely would also transcend. Instead, as predicted, Chessformer 1500 likely is not transcendent due to a lack of diversity in its dataset.

### v2

@@ Abstract
Generative models are trained with the simple objective of imitating the conditional probability distribution induced by the data they are trained on. Therefore, when trained on data generated by humans, we may not expect the artificial model to outperform the humans on their original objectives. In this work, we study the phenomenon of transcendence: when a generative model achieves capabilities that surpass the abilities of the experts generating its data. We demonstrate transcendence by training an autoregressive transformer to play chess from game transcripts, and show that the trained model can sometimes achieve better performance than all players in the dataset. [^1] To play with our models, code, and data, please see our website at https://transcendence.eddie.win.  We theoretically prove that transcendence can be enabled by low-temperature sampling, and rigorously assess this claim experimentally. Finally, we discuss other sources of transcendence, laying the groundwork for future investigation of this phenomenon in a broader setting.
@@ 1 Introduction
Our objective is to formalize the notion of transcendence and focus narrowly on this source of improvement over the experts: the removal of diverse human biases and errors. We prove that this form of denoising is enabled by low-temperature sampling, which implicitly induces a majority vote. Our result draws a subtle but deep connection from our new setting to a rich prior literature on model ensembling [1, 6, 19], enabling several key results. We precisely characterize the conditions under which transcendence is possible, and give a rigorous theoretical framework for enabling future study into the phenomenon. To test the predictive power of our theory, we then empirically demonstrate these effects. Digging deeper into the effects of majority voting, we show that its advantage is primarily due to performing much better on a small subset of states—that is, under conditions that are likely key to determining the outcome of the game. We also find that diversity in the data is a necessary condition for practically effective majority voting, confirming our theoretical findings. In short:
@@ 4.1 Experimental Setup
We trained several $50$M parameter autoregressive transformer decoders following best practices from modern large model training, including a cosine learning rate schedule and similar batch size-learning rate ratios as prescribed by the OPT-175B team [37]. Our dataset consists of human chess games from the lichess.org open source database from January 2023 to October 2023. In total, this dataset contains approximately one billion games. In this setting, an expert is a specific individual player. To test for transcendence, we truncate this dataset by a maximum rating, so that during training a model only sees data up to a given rating. We train our model on the next-token prediction objective, and represent our chess games as Portable Game Notation (PGN) strings, such as 1.e4 e5 2.Nf3 Nc6 3.Bb5... 1/2-1/2. Note that we do not give any rating or reward information during training—the only input the model sees are the moves and the outcome of the game. We tokenize our dataset at the $32$-symbol character level. (For further details, see Appendix E.) Our model plays chess “blind”—without direct access to the board state—and, furthermore, is never explicitly given the rules of the game: at no point is play constrained to valid outputs for a given piece or board state. Nontrivial chess skill is therefore not straightforward to acquire, and if not for the surprising capabilities of modern large transformers, one might imagine such a model would fail to learn even the basic rules of playing chess. This blindfolded setting has also been studied by prior work [23, 30], as discussed further in section 5.
One gap between our theory and practice is that in our theory, we assume that each expert is defined over the entire input space $\mathcal{X}$. However, in the chess setting such full coverage is extremely unlikely to be the case after around move $15$, as there are more unique chess games than atoms in the universe due to the high branching factor of the game tree. To address this gap, we visualize the latent representation of our model in Figure 3, where we find the model is able to capture meaningful semantics regarding both the relative advantage of a state, as well as the identity of the black and white player. This visualization illustrates the ability of our model to generalize by compressing games into some shared latent representation, enabling experts to generalize to unseen states, bridging this gap between theory and practice.
@@ 4.2 Experimental Results
Where $d^{f}$ refers to the state visitation distribution [31] when following $f$ in a sequential setting—informally, this variable can be thought of the distribution of states seen when sampling from $f$ with a fixed transition function that takes in an input $x$, a output $y$, and outputs a next input $x$. Here, that transition function is given by the rules of chess and the opponent player. Given this favor function, we can now quantitatively explore the effects that lead to transcendence by setting the baseline $f$ to be the original imitation-learned probability distribution (temperature $\tau=1$), and $f^{\prime}$ as a low-temperature intervention on $f$ (e.g. temperature $\tau=0$). We can empirically calculate the reward by using the evaluation function [22] of Stockfish, an expert neural reward function that Stockfish uses to calculate its next move. This reward function is a neural network trained to predict the probability of winning through a sigmoid on a linear combination of handcrafted expert heuristics, such as amount of material versus opponent material, and number of moves to a potential checkmate.
@@
In Figure 4, we find that lowering the temperature has the effect of skewing the expected reward distribution to the left, especially for the green $\tau=0.001$ distribution. This result implies that the model does not improve the expected reward by a small amount for many game states, but rather improves the expected reward by a relatively large amount for a few game states. Thus, $\tau=0.001$ improves the expected reward (probability of winning) by an average of $\mathbf{2.15\pm 0.17\%}$, but for some states, this expected improvement is over 5%. Note that the original temperature expected reward can be thought of as a Dirac distribution centered at 00.  The above finding answers our research question in this section: Low-temperature sampling is able improves the expected reward by relatively large amounts for some specific game states, which is likely why the ChessFormer $1000$ and $1300$ model was able to achieve transcendence.
@@
In Table 1, we present the statistics of the favor function for different temperature values. From this table, we observe that as the temperature decreases, the top-$k$ accuracies monotonically increase, suggesting that the model becomes more consistent in selecting good moves. We also observe that although the model improves as temperature decreases, the probability of winning is still below $50\%$, meaning our model should tend to lose more games than it wins against Stockfish $1$. This result matches with our results in Figure 1, as the rating of Stockfish $1$ is also higher than the reported rating for $\tau=0.001$ ($1550$ for Stockfish 1 vs $\sim 1450$ for Chessformer $1000$). Overall, the analysis of the advantage statistics provides further evidence for the effectiveness of low-temperature sampling in inducing transcendence in chess models.
@@
As we note in subsection 3.4, our theory requires dataset diversity as a necessary condition for enabling transcendence. Importantly, we find in Figure 1 that not all models are able to transcend. Unlike ChessFormer 1000 or 1300, the Chessformer 1500 fails to transcend. We hypothesize that this results is due to the fact that in the band of ratings from $1000$ to $1500$, diversity does not significantly increase. If so, a $1000$ rated player can be thought of as a noisy $1500$ rated player, but a $1500$ rated player cannot be thought of as a noisy $2000$ rated player. In this section we ask the following research question: Is diversity in data required for enabling transcendence?
@@
Importantly, we cannot calculate this normalized entropy for every state, as most states after move $16$ in the midgame and before the engame are unique within the dataset and we therefore observe just a single action for thus states. Therefore our metric is limited in that it only considers opening moves, the beginning of the midgame, and the endgame. We consider only common states with greater than $100$ actions by sampling $1,000,000$ games from each dataset. The average entropy confirm our hypothesis: The $<1500$ cut off dataset has on average less diversity than the $<1300$ dataset, which has is again less than the $<1000$ dataset. This result suggests that Chessformer $1500$ likely is not transcendent due to a lack of diversity in its dataset. If the entropy instead stayed constant for each dataset, it would imply that each had a similar level of diversity. In such a case, we would expect that ChessFormer $1500$ likely would also transcend. Instead, as predicted, Chessformer 1500 likely is not transcendent due to a lack of diversity in its dataset.
@@ 6 Discussion and Future Work
Acknowledgements
Sham Kakade acknowledges this work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence; support from the Office of Naval Research under award N00014-22-1-2377, and the National Science Foundation Grant under award #IIS 2229881.